Nora, whose mass is 50.0 kg, stays out skiing for too long and her body temperature drops by 2.00°C. What is the amount of heat lost from Nova’s body? (chuman body = 3470 J/kg°C)
1
Expert's answer
2013-12-14T08:17:26-0500
Q=cm(t2-t1)=3470*50*2=347000 J = 347 kJ